# Scalable Vision Language Model Training via High Quality Data Curation

### v1

@@ 3.1 Scalable High-Quality Visual Understanding Data Construction
To quantitatively demonstrate the advantages of the SAIL-Caption dataset over other open-source datasets, we curated a 500-case subset from each dataset for data quality evaluation. The detailed evaluation process is elaborated in Appendix B.1. Results show that SAIL-Caption-DataComp and SAIL-Caption-SA1B outperform previous baseline datasets by $20.35\%$ and $11.91\%$, respectively, exhibiting considerable data quality improvements. We also show visualized showcases of SAIL-Caption in Appendix B.2.
Figure 2: Scaling curves of SAIL-VL’s performance dynamics in pretrain-advance stage. We show model performance on all understanding benchmarks, caption tasks and OCR tasks, respectively.
Figure 3: Scaling curves of SAIL-VL’s performance dynamics in pretrain-advance stage. We show average model performance on all VLM benchmarks used in our experiments, general VQA tasks, OCR VQA tasks, respectively. The two sub-domains cover visual instruction following tasks for natural scenes and complicated documents, respectively. The detailed benchmark split for different domains is shown in Table 2.
In this part, we introduce the scaling laws observed in SAIL-VL large-scale pretraining. We scale up data size used in SAIL-VL pretrain-advance stage, with training budgets configured accordingly. For model checkpoints obtained at different pretraining steps, we conduct a lightweight annealing training to let model parameters converge to local optimal points, ensuring the stability of evaluation results.
SAIL-VL is trained through 16B tokens for initial vision-language alignment, and then ingests 131B tokens during the pretrain-advance stage, during which we investigate model performance dynamics. To evaluate the visual understanding performance of SAIL-VL, we establish an evaluation suite which covers fundamental visual understanding tasks such as detail caption generation Dong et al. (2024) and OCR detection Biten et al. (2022); Wang et al. (2020); Gupta et al. (2016); Kim et al. (2022). Details can be found in Appendix D.
We select checkpoints trained with exponentially larger data sizes from the pretrain-advance stage to inspect model performance dynamics. As shown in Figure 2, SAIL-VL’s overall performance, captioning performance, and OCR detection performance improve smoothly as the training proceeds. As pretraining data sizes and training budgets scale up exponentially, SAIL-VL’s visual understanding performance improves linearly. This observation unveils a promising prospect to scale up VLM pretraining data sizes for improved model performance.
To further investigate the effectiveness of SAIL-VL’s large-scale pretraining, we conduct visual instruction tuning with model checkpoints collected from the pretrain-advance stage. Specifically, we curate visual instruction tuning data from the SFT-instruction stage, and evaluate the resulting model checkpoints on 19 commonly used VLM benchmarks to validate the scaling results.
As shown in Figure 3, SAIL-VL’s overall performance improves steadily as the data size and computation budget increase, and the dynamics can be plotted as a near-linear curve on an exponential horizontal axis. Even at the data point derived from the model checkpoint trained with 131B tokens, the scaling curve of SAIL-VL’s performance does not converge. We also show the scaling results on two widely used benchmark series: General VQA and OCR VQA. As the training budget scales up exponentially, SAIL-VL’s consistently develops more robust abilities to recognize visual elements and tackle instruction following tasks, leading to near-linear performance scaling curves.
3.3 Scaling up VLM Instruction Tuning
The instruction tuning (SFT) stage is crucial for VLMs’ downstream performance. While the visual instruction tuning data from the internet is abundant, high-quality data is still scarce. To solve this problem, we introduce several guidelines for visual instruction tuning data curation in Section 3.3.1, and demonstrate model performance scaling laws of the curriculum SFT strategy with progressively improved data quality in Section 3.3.2.
3.3.1 Instruction Data Curation for Quantity Scaling
In this section, we elaborate on the guidelines for visual instruction tuning data curation and demonstrate their effectiveness by training with our customized SFT data.
Efficient data quality evaluation.
Firstly, we propose an intuitive way for data quality evaluation. For any dataset, we train a model on a small subset of data with a fixed quantity (2M samples) for efficiency, and then conduct holistic evaluation on a series of benchmarks. The evaluation results reflect training data quality comprehensively. The SFT data distribution is decided with reference to the data quality evaluation results.
Detecting low-quality components from SFT data mixing.
To detect which part of visual instruction data undermines model performance the most, we propose Composition Evaluation, which identifies data quality of SFT data components from model performance variations by halving each part of the data. We conduct composition evaluation on each SFT data category, and repeat this process until no further improvement is observed. Equipped with the proposed composition evaluation strategy, we improve SAIL-VL’s SFT-instruction training data distribution across existing SFT data sources to achieve improved model performance.
Figure 4: Scaling curves of model performance trained on our SAIL-Instruct dataset, LLaVA-OneVision Li et al. (2024a) single image SFT data, and datasets from Infinity-MM Gu et al. (2024).
Identifying incoming visual instruction tuning data quality.
For incoming visual instruction tuning data to be incorporated, we explore scaling up SFT data sizes without compromising overall data quality. We propose the Incremental Evaluation strategy, which identifies the quality of an incoming visual instruction tuning dataset from model performance variations with the incoming dataset included in the training data. This simple-yet-effective approach enables us to continuously expand data quantity while preserving its quality.
Figure 5: Model performance dynamics of the quality scaling and all-in-one (AIO) training strategy. “AIO learning” incorporates all three-stage SFT data into a single training loop.
Quantity Scaling.
Based on this methodology, we collect a large amount of high-quality visual instruction tuning data. As shown in Figure 4, model performance scales up stably as the data quantity grows from 2M to 12M, depicting a logarithmic scaling curve as training data quantity increases. Training with datasets of different quality results in varying model performance. Among these datasets, our SAIL-Instruct dataset achieves the best quality evaluation results, surpassing recently proposed alternatives either at the beginning of model training or at the end of the scaling curves.
3.3.2 Multi-stage Instruction Tuning with Quality Scaling
In this part, we introduce quality scaling, a curriculum learning strategy for VLM SFT, which starts with abundant low-quality data and switches to scarce high-quality data at the end of the training, to achieve even higher model performance than one-stage SFT.
Curriculum SFT with progressively improving data quality.
We train SAIL-VL through three SFT stages in a curriculum learning paradigm. SAIL-VL ingests world knowledge and develops basic visual instruction following abilities during the first two stages, and learns to tackle complex tasks in the last training stage. As elaborated in Section 2.2, visual instruction tuning data used in the three SFT stages differs in the following ways: (1) Datasets with more scarce data quantity are used for later stages. (2) Data formats become increasingly complex across the stages. (3) Data quality progressively improves through the training process. The three-stage model training strategy ensures effective utilization of scarce-yet-high-quality datasets, and therefore enables SAIL-VL to develop robust and comprehensive visual instruction following abilities.
Quality Scaling.
We show model performance dynamics derived from the three SFT stages in comparison with an all-in-one (AIO) training strategy in Figure 5. The model trained with the curriculum SFT strategy exhibits a near-linear performance scaling curve across training stages with progressively improved data quality, yielding better final model performance over AIO training. Model performance dynamics during AIO training can be described as a logarithmic curve with suboptimal performance. We attribute this performance inferiority to the inefficient data allocation strategy compared to our curriculum SFT approach.
@@ 4 Experiments
Benchmark SAIL-VL Qwen2-VL InternVL2 InternVL2.5 Aquila DeepSeekVL-2-Tiny
Overall Performance
Opensource Average 68.6 64.0 62.3 66.1 65.9 66.8
   General VQA 60.2 58.2 55.3 62.1 59.2 59.9
   OCR VQA 75.7 72.5 68.2 68.7 71.7 74.4
   Math&Knowledge 79.6 59.0 70.6 73.0 75.4 71.3
   Hallucination 66.9 62.9 61.7 66.2 62.9 63.6
OpenCompass Average 61.3 55.2 54.1 59.6 58.7 57.3
General VQA
MMStar (2024b) 56.6 46.3 50.5 53.5 54.7 49.9
MMBench (2024b) 74.4 68.8 70.3 73.1 74.4 68.3
MMMU (2024b) 42.0 39.9 34.2 40.7 44.1 39.6
MME (2024b) 2001 1923 1860 2091 1808 1911
SEEDBench (2024b) 74.7 72.0 70.9 73.4 73.9 72.5
RealWorldQA (2024b) 62.7 60.9 56.7 60.9 64.1 64.8
MMVet (2023) 47.0 51.2 40.4 61.7 42.7 52.8
LLaVABench (2024b) 52.3 57.5 52.8 59.0 54.8 62.8
OCR VQA
AI2D (2024b) 77.3 72.3 74.2 75.0 75.0 74.6
DocVQA (2024b) 88.3 88.7 86.0 87.4 85.0 88.6
InfoVQA (2024b) 66.5 63.4 57.5 61.6 60.5 63.8
ChartQA (2024b) 79.6 70.6 71.7 73.3 76.6 81.2
TextVQA (2024b) 77.3 78.8 73.4 76.4 76.4 80.5
OCRVQA (2024b) 57.2 54.3 36.2 28.3 51.3 51.4
OCRBench (2024b) 835 794 786 789 772 808
Math&Knowledge
MathVista (2024b) 63.4 45.0 46.8 51.1 59.4 54.5
ScienceQA (2024b) 95.8 73.0 94.4 94.9 91.4 88.1
Hallucination
HallusionBench (2024b) 46.0 38.3 38.2 42.5 42.1 38.4
POPE (2024b) 87.8 87.6 85.3 89.9 83.6 88.8
Table 2: Evaluation results of SAIL-VL and other opensource VLM with comparable sizes. “Opencompass Average” indicates average model performance on OpenCompass benchmarks, while “Opensource average” includes all opensource benchmarks listed in the table. Bold numbers indicate the best performance on the benchmark, while underlined ones are those ranked as the second.
@@ 4.1 Experiment Setup
Training framework.
Training resources.
We conduct experiments with Huawei 910B x86 NPU. A total of 26,656 NPU hours are consumed to train SAIL-VL-2B, with pretraining stages and SFT stages consuming 15,564 and 10,992 NPU hours, respectively.
@@
We compare our SAIL-VL-2B model with previous SOTA VLM baselines of comparable sizes, including Qwen2-VL-2B Wang et al. (2024), InternVL2-2B Team (2024a), InternVL2.5-2B Chen et al. (2024c), Aquila-2B Gu et al. (2024), and DeepSeekVL-2-2BWu et al. (2024).
@@
We evaluate SAIL-VL and baseline VLMs on a series of widely used benchmarks. We categorize the opensource benchmarks used in our evaluation pipeline into four types: General VQA, OCR VQA, Math&Knowledge, and Hallucination, covering VQA tasks on natural images/videos, OCR-related documents, as well as those involving complicated reasoning and world knowledge to tackle. All results are evaluated in our customized version of VLMEvalKit Duan et al. (2024), which may differ slightly from model performance reported in other papers or the official OpenCompass leaderboard.
@@ 4.2 Benchmark Results
SAIL-VL ourperforms previous SOTA VLMs with comparable sizes significantly.
We list the performance of SAIL-VL along with other opensource VLMs in Table 2. As the results show, SAIL-VL outperforms previous SOTA VLMs by a large margin w.r.t. average scores across all benchmarks or the OpenCompass benchmark collection, indicating robust visual understanding and instruction following abilities. SAIL-VL achieves the top-2 performance on 13 out of 19 benchmarks in our evaluation. Compared to previous SOTA VLMs of comparable sizes, SAIL-VL scores $1.8$ ($2.69\%\uparrow$) higher opensource average performance than DeepSeekVL-2-Tiny, and $1.7$ ($2.85\%\uparrow$) higher OpenCompass average score than InternVL-2.5, respectively.
SAIl-VL scores the highest in visual comprehension across in almost every domain.
It is worth noticing that SAIL-VL achieves leading performance across benchmarks of various domains. SAIL-VL exceeds previous SOTA VLMs by $1.3$ ($1.75\%$) in OCR VQA, $4.2$ ($5.57\%$) in math&knowledge tasks, and $0.7$ ($1.06\%$) in Hallucination benchmarks, respectively. SAIL-VL’s performance decays slightly in general VQA tasks, especially in the MMVet benchmark. We attribute this performance drop to the open-ended answering paradigm, which may introduce instability in evaluation because of the varying answering lengths of different VLMs. These results indicate that SAIL-VL develops powerful visual comprehension and language instruction abilities within its limited model size, demonstrating promising prospects in downstream applications.
@@ 5 Analysis
5.1 Training Pipeline Ablation
PT-adv SFT-knw SFT-inst SFT-prf OS Avg
68.6
67.7
67.7
66.2
67.1
Table 3: SAIL-VL’s performance with each training stage removed. ✓and ✗ indicate that this stage is retained and removed, respectively. “OS Avg” stands for average model performance on the 19 opensource benchmarks.
We conduct ablation experiments to show the necessity and effectiveness of our five-stage training pipeline. We experiment with each training stage removed except the pretrain-alignment stage, because a randomly initialized projector may cause ViT parameter distribution collapse during the subsequent training stages. Experiment results are shown in Table 3, where a consistent performance drop is witnessed with any training stage removed, validating the necessity and effectiveness of our training pipeline design.
@@ 5.2 Pretrain Data Quality Determines Pretrained Model Performance
We explore pretraining SAIL-VL with caption data of varying data quality. Specifically, we conduct lightweight experiments in the pretrain-advance stage, starting from the model checkpoint after the same alignment pretraining. A total of 16B tokens are consumed for these ablation studies, consisting of a fixed amount of OCR and caption data from different data sources. We evaluate model performance with our understanding benchmarks.
Caption Data Understanding Avg. Caption Avg. OCR Avg.
SAIL-Caption 52.37 51.19 53.54
SA1B-QwenVL-Caption 50.97 48.20 53.74
DataComp-LLaVA-Caption 48.90 44.92 52.88
Table 4: Visual understanding performance of model checkpoints pretrained with different data sources. We report models’ average visual understanding performance on caption and OCR benchmarks, respectively.
@@
As shown in Table 4, the model trained with SAIL-Caption achieves significantly higher performance than those trained on other opensource caption datasets. The model performance evaluation results coincide smoothly with our data quality assessment findings: the higher the pretrain caption data quality, the better the pretrained model performs. Details about our data quality evaluation results are shown in Appendix 6.
5.3 Generalizing Visual Understanding Abilities to Instruction Following Tasks
Figure 6: The correlation between SAIL-VL pretrained checkpoints’ understanding performance and their performance on opensource benchmarks after SFT. “OS BMK Score” stands for average score on opensource benchmarks.
In this part, we investigate the correlation between VLM’s visual understanding and instruction following abilities. We train SAIL-VL pretrained checkpoints from the pretrain-advance stage with exponentially larger training data sizes, and train them through either LLaVA-Next Liu et al. (2024a) SFT data or our SFT-instruction data. We plot the correlation of pretrained models’ visual understanding performance and SFT models’ opensource benchmark performance in Figure 6. A notable correlation is observed across different training strategies. As the VLM gains stronger visual understanding abilities during pretraining, its visual instruction following performance after SFT is improved accordingly, even if trained with different visual instruction tuning datasets. We quantify this correlation with Pearson correlation ($\rho$) and coefficient of determination ($R^{2}$). It turns out that SAIL-VL’s pretrained visual understanding performance and SFT visual instruction following performance share a significant correlation with $\rho=0.91$ and $R^{2}=0.83$, validating the generalization of model abilities across training stages.
5.4 Data Composition and Quality Evaluation for Visual Instruction Tuning
Data Composition Quality Evaluation
Stages Size General OCR Math. Text Overall General OCR Math. Hall.
Knowledge Learning 21M 36% 43% 7% 14% 0.548 0.511 0.582 0.556 0.568
Instruction Learning 12M 44% 31% 9% 16% 0.609 0.542 0.657 0.681 0.631
Preference Learning 3.5M 76% 11% 10% 3% 0.606 0.560 0.658 0.595 0.613
@@ 5.4 Data Composition and Quality Evaluation for Visual Instruction Tuning
Table 5: Composition and quality evaluation of dataset from each stage. Note that Math. for data composition includes data with math and reasoning, Math. for evaluation denotes Math &Knowledge, Hall. denotes Hallucination.
To demonstrate the differences between SFT stages, we show the data distribution of each stage along with the performance of models trained with each part of SFT data. As illustrated in Table 5, the composition of knowledge learning and instruction learning data is similar, but the resulting model performance improves significantly. The composition of preference learning data differs from data used in other stages, containing much more open question-answering samples. Although training with preference learning data does not yield the highest performance scores, model performance in general VQA tasks. With a small amount of high-quality OCR samples, the OCR evaluation score for preference learning is also high.
@@ 6.1 Visual Understanding Data
Visual understanding data consists of vision modality contents and corresponding language depictions, and is regarded as the keystone to various vision and language model applications. Whether it is representation learning models like CLIP and its derivatives Radford et al. (2021); Jia et al. (2021); Shen et al. (2022); Cherti et al. (2023); Fang et al. (2023), generative models Wang et al. (2022); Li et al. (2021); Bao et al. (2022); Yu et al. (2022); Li et al. (2023), or recent vision language models Li et al. (2023); Liu et al. (2024b); Team (2024a); Bai et al. (2023), all of these methods are built upon large scale high-quality visual understanding data. LAION Schuhmann et al. (2021, 2022), TaiSu Liu et al. (2022), Coyo Byeon et al. (2022), and etc. provide relatively low-quality alt-texts paired with source images. Subsequent works such as CapsFusion Yu et al. (2024), SA1B-Recaption Data (2024b), DataComp-Recaption Li et al. (2024b), and etc. explore enhancing the caption data quality with recaption techniques, and the resulting datasets are widely used in recent VLM researches.

### v3

@@ 3.3.2 Multi-stage Instruction Tuning for Data Complexity Scaling
$\text{DocVQA}_{\text{VAL}}$ (2021) 89.2 88.7 87.8 88.6 92.2 93.7 92.1 92.3
$\text{InfoVQA}_{\text{VAL}}$ (2022) 67.2 63.4 61.6 63.8 75.2 75.9 76.2 72.5
$\text{ChartQA}_{\text{TEST}}$ (2022) 81.0 70.6 70.9 81.2 84.6 81.6 77.6 84.6
$\text{TextVQA}_{\text{VAL}}$ (2019) 75.7 78.8 77.2 80.5 77.7 83.8 79.2 83.3
$\text{OCRVQA}_{\text{TEST}}$ (2019) 58.5 54.3 40.0 51.4 61.4 56.2 36.7 54.5
OCRBench (2024d) 806 794 846 808 835 833 880 834
Math&Knowledge
$\text{MathVista}_{\text{MINI}}$ (2023) 62.8 45.0 55.3 54.5 68.4 57.3 68.5 61.8
$\text{ScienceQA}_{\text{VAL}}$ (2022) 95.3 73.0 95.3 88.1 98.2 84.6 97.9 96.2
Hallucination
HallusionBench (2024) 45.7 38.3 39.2 38.4 52.2 48.5 50.3 41.2
POPE (2023) 86.7 87.6 89.8 88.8 85.2 86.5 89.1 89.4
Table 4: Evaluation results of SAIL-VL and other opensource VLM with comparable sizes. “Opensource average” includes all opensource benchmarks listed in the table. Bold numbers indicate the best performance among models of comparable sizes, while underlined ones are those ranked as the second.
@@ 4.1 Experiment Setup
Model Training.
We start from InternViT-300M Chen et al. (2023b), Qwen2.5-2B and Qwen2.5-7B Team (2024b) for model training. Detailed model training recipes are elaborated in Table 3. As training progresses, the input image resolution gradually increases, with a $2\times 2$ pixel shuffle Chen et al. (2024d) module employed in the projector, maintaining a balance between efficiency and performance. For SAIL-VL-8B, we use smaller batch sizes and larger learning rates in pretraining stages to improve training efficiency. During SFT stages, the 8B model is trained with a smaller learning rate, mitigating the instability in full model training with larger LLMs.
@@
We compare our SAIL-VL models with previous SOTA VLM baselines of comparable sizes, including Qwen2-VL Wang et al. (2024b), InternVL2.5-MPO Chen et al. (2024c), DeepSeekVL-2Wu et al. (2024), etc. Evaluation results against more existing baseline models are shown in Appendix H.2.
@@
We evaluate SAIL-VL and baseline VLMs on a series of widely used benchmarks, including General VQA, OCR VQA, Math&Knowledge, and Hallucination. These categories cover VQA tasks on natural images/videos, OCR-related documents, as well as those involving complicated reasoning abilities and world knowledge to tackle. We use a customized version of VLMEvalKit Duan et al. (2024) for evaluations.
@@ 4.2 Benchmark Results
SAIL-VL-2B ourperforms previous SOTA VLMs with comparable sizes significantly.
We list the performance of SAIL-VL along with other opensource VLMs in Table 4. As the results show, SAIL-VL-2B outperforms previous SOTA VLMs by a large margin, scoring $1.4$ ($2.06\%\uparrow$) higher average performance than InternVL2.5-MPO-2B. SAIL-VL-2B achieves new SOTA performance on 3 out of 4 subfields except for General VQA. We attribute it to the instability lying in benchmarks requiring long text generation, such as MMVet.
SAIl-VL-8B achieves leading performance over opensource baselines.
As shown in Table 4, SAIL-VL-8B also achieves leading visual comprehension performance over Qwen2-VL, DeepSeekVL-2, and even InternVL2.5-MPO-8B, which requires an additional reinforcement learning stage in model training. We admit the shrunk performance advantage of SAIL-VL-8B over SOTA baselines, which may be caused by the relatively small data sizes used for model training. We take these results as an early attempt for larger VLM training, and more competitive large VLMs will be released in our SAIL-VL series.
@@ 5.1 Pretrain Data Quality Determines Pretrained Model Performance
We explore pretraining SAIL-VL-2B with varying data quality. Specifically, we conduct lightweight 16B-token training in the pretrain-advance stage, starting from the model checkpoint after the same alignment pretraining. We fix the data distribution across different data types, and modify data composition with varying-quality data.
@@
As shown in Table 5, the model trained with SAIL-Caption achieves significantly higher performance than those trained on other opensource caption datasets, which is consistent with data quality evaluation results as shown in Appendix 8. It is also worth noticing that the model trained with repeated-yet-high-quality OCR data yields better results than incorporating diverse but relatively low-quality data for model training. We attribute this result to our frozen-LLM pretraining setting, which mitigates the potential overfitting problem lying in repeated training data.
Caption Data OCR Data Overall Caption OCR
SAIL-Caption HQ 54.36 51.80 55.38
SA1B-QwenVL-Caption HQ 48.43 39.57 51.97
DataComp-LLaVA-Caption HQ 49.08 42.70 51.63
BLIP3-KALE HQ 53.06 46.00 55.89
SAIL-Caption HQ+LQ 52.13 51.22 52.50
SAIL-Caption HQ (RP) 54.05 52.63 54.62
Table 5: Visual understanding performance of model checkpoints pretrained with different data sources. We report models performance on our visual understanding benchmarks. “HQ”, “LQ”, and “RP” indicates high-quality, low-quality, and repeated data, respectively.
5.2 SFT Data Quality Analysis
To further validate our data quality evaluation results shown in Table 1, we select 2M-sample subsets from each SFT stage to train the pretrained SAIL-VL-2B model. Performance evaluation results are shown in Table 6. A significant performance advantage is observed in the model trained with SFT-Instruction data collection, validating the effectiveness of the proposed data curation methods. This result coincides with the data quality evaluation results given in Table 1, where SFT-Instruction data collection exhibits advanced task difficulty, data complexity, and image-text relevance. It is also worth noticing that despite the improved data quality of the SFT-Preference data, it fails to further improve model performance in Table 6. We attribute it to its excessively high data complexity, which may hinder effective model learning. This observation further validates the proposed curriculum VLM SFT strategy as discussed in Section 3.3.2.
Training Data Overall General OCR Math. Hall.
SFT-Knowledge 57.8 53.2 60.9 56.9 63.9
SFT-Instruction 61.9 55.8 67.1 65.4 61.7
SFT-Preference 61.3 57.1 65.8 59.5 61.3
@@ 5.2 SFT Data Quality Analysis
Table 6: Performance evaluation results of models trained with SFT data from each stage. We denote “Math.” as Math &Knowledge benchmarks in evaluation. “Hall.” denotes Hallucination benchmarks as defined in Section 4.1.
@@ 6.1 Visual Understanding Data
Visual understanding data consists of vision modality contents and corresponding language depictions, and is regarded as the keystone to various vision and language model applications. Whether it is representation learning models like CLIP and its derivatives Radford et al. (2021); Jia et al. (2021); Shen et al. (2022); Cherti et al. (2023); Fang et al. (2023), generative models Wang et al. (2022); Li et al. (2021); Bao et al. (2022); Yu et al. (2022); Li et al. (2023b), or recent vision language models Li et al. (2023b); Liu et al. (2024b); Team (2024a); Bai et al. (2023), all of these methods are built upon large scale high-quality visual understanding data. LAION Schuhmann et al. (2021, 2022), TaiSu Liu et al. (2022), Coyo Byeon et al. (2022), DataComp Gadre et al. (2024), and etc. provide relatively low-quality alt-texts paired with source images. Subsequent works such as ShareGPT4V chen2023sharegpt4v and ALLaVA Chen et al. (2024a) annotate small scale high-quality caption data with powerful VLM APIs. To produce high-quality detail caption data at scale, CapsFusion Yu et al. (2024a), World2Seq Wang et al. (2024a), CAPTURE Dong et al. (2024), SA1B-Recaption Data (2024b), DataComp-Recaption Li et al. (2024b), and BLIP3-KALE Awadalla et al. (2024) employ recaptioner models for efficient data annotation. The resulting datasets are widely used in recent VLM research.